It took Cindy 2 hours to bike from Abbott to Benson at a constant speed. The return trip took only 1.5 hours because she increased her speed by 6 mph. How far apart are Abott and Benson? How fast did go to get to Benson? How fast did she return?
Answer by Paul(988)

Let the current speed be x
Let the return speed be x+6

1.5(x+6)=2x
1.5x+9=2x
0.5x=9
x=18

18+6=24

1.5(18+6)=2(18)
36=36

Hence, she went at 18mph and returned at 24mph and the total distance between is 36 miles.
Paul.